OpenStudy (anonymous):

0.2x+5=22 WILL GIVE MEDAL FOR CORRECT ANSWER!

jimthompson5910 (jim_thompson5910):

0.2x+5=22 0.2x=22-5 0.2x = 17 x = ???

OpenStudy (anonymous):

85?

jimthompson5910 (jim_thompson5910):

yep